Determmine value of k such that



Kx+y+z=1



X+ky+z=1



X+y+kz=1 has a) no solution b) unique solution c) more than one solution





1
Expert's answer
2022-02-28T15:32:55-0500
"\det A=\begin{vmatrix}
 k & 1 & 1 \\
 1 & k & 1 \\
 1 & 1 & k
\end{vmatrix}=k\begin{vmatrix}
 k & 1 \\
 1 & k
\end{vmatrix}-\begin{vmatrix}
 1& 1 \\
 1 & k
\end{vmatrix}+\begin{vmatrix}
 1 & k \\
 1 & 1
\end{vmatrix}"

"=k(k^2-1)-(k-1)+(1-k)"

"k^3-k-k+1+1-k=k^3-3k+2"

"=k^2(k-1)+k(k-1)-2(k-1)"




"=(k-1)(k^2+k-2)=(k-1)^2(k+2)"

Nonhomogeneous system of linear equations has a unique non-trivial solution if and only if


"\det A\not=0=>(k-1)^2(k+2)\not=0"

The system has a unique non-trivial solution if "k\in \R, k\not=1, k\not=-2."


If "k=1," we have


"\begin{matrix}
 x+y+z=1\\
 x+y+z=1\\
 x+y+z=1\\
\end{matrix}"

The system has an infinite number of solutions if "k=1."


If "k=-2," we have


"\begin{matrix}
 -2x+y+z=1\\
 x-2y+z=1\\
 x+y-2z=1\\
\end{matrix}"




"\begin{matrix}
 -3x+3y=0\\
 3y-3z=0\\
 x+y-2z=1\\
\end{matrix}"




"\begin{matrix}
 x=y\\
 y=z\\
 y+y-2y=1\\
\end{matrix}"



The system has no solution if "k=-2."